OpenStudy (anonymous):

As Juliet stands 15 feet off the ground on her balcony, she spots Romeo in this distance, standing at the ground level. She uses her angle-measuring device and finds that his position is 3 degree from the horizontal (that is, the angle of depression is 3 degree.) To the nearest foot, how far away is Romeo from base of her balcony?

OpenStudy (anonymous):

The angle to the ground is 90-3=87 this gives us the opposite and adjacent which is tan so \[\tan(87) =\frac{opp}{15} = 15 \times \tan (87) = distance\] distance is 286.217 to the nearest foot is 286 feet